# Theoretical Bounds for Stable In-Context Learning

### v1

@@ Abstract
In-context learning (ICL) is flexible but its reliability is highly sensitive to prompt length. This paper establishes a non-asymptotic lower bound that links the minimal number of demonstrations to ICL stability under fixed high-dimensional sub-Gaussian representations. The bound gives explicit sufficient conditions in terms of spectral properties of the covariance, providing a computable criterion for practice. Building on this analysis, we propose a two-stage observable estimator with a one-shot calibration that produces practitioner-ready prompt-length estimates without distributional priors. Experiments across diverse datasets, encoders, and generators show close alignment between the predicted thresholds and empirical knee-points, with the theory acting as a conservative but reliable upper bound; the calibrated variant further tightens this gap. These results connect spectral coverage to stable ICL, bridge theory and deployment, and improve the interpretability and reliability of large-scale prompting in realistic finite-sample regimes.
@@ 1 Introduction
In-Context Learning (ICL) has become a central capability of large language models, enabling them to solve novel tasks at test time without parameter updates. Despite its flexibility, the reliability of ICL is highly sensitive to the number of demonstrations provided in the prompt. Practitioners report that the stability threshold varies widely across tasks and models, and naive scaling rules often fail to predict it accurately [1, 2]. This motivates a principled question: what is the minimal number of demonstrations required for stable ICL?
Recent theory provides partial answers. PAC-style analyses study when ICL is learnable and give finite-sample guarantees under mixture assumptions [3]. Algorithmic perspectives view transformers as learners with task-dependent sample complexity, yielding upper bounds rather than minimal requirements [4]. Asymptotic learning curves characterize phase transitions in analytically tractable models [5]. Complementary lines of work improve calibration in ICL [6, 7, 8], or reveal mechanistic insights into emergent behavior [9, 10, 11]. However, none of these studies provide a non-asymptotic, computable lower bound that directly links prompt length to stability via spectral properties [12, 13].
In this work, we address this gap by theoretically analyzing the relationship between the number of demonstrations in a prompt and the stability of ICL. Assuming the model relies on fixed high-dimensional representations, we derive an explicit non-asymptotic lower bound that characterizes when ICL remains stable with high probability. Building on this analysis, we propose a two-stage observable method with one-shot calibration that operationalizes the result in practice without requiring distributional priors [14].
Our contributions are threefold:
  • We provide a non-asymptotic theoretical lower bound that, for the first time, establishes a quantitative relationship between the number of in-context demonstrations and the stability of ICL [3].
  • We design an observable two-stage algorithm that estimates the required number of demonstrations in practical tasks while preserving theoretical guarantees [4, 14].
  • We conduct extensive experiments to validate the consistency between our theoretical predictions and empirical model behavior [1, 2].
@@ 2 Related Works
Theoretical analyses.
Prior studies have examined ICL from different theoretical angles. PAC-style formulations establish when ICL is learnable under mixture models [3]. Transformers have also been analyzed as algorithms with task-dependent sample complexity [4], while asymptotic learning curves describe phase transitions in analytically tractable settings [5]. Task-specific results for regression and structured prediction further refine these analyses [15].
Calibration and reliability.
Calibration-based methods improve few-shot prediction confidence and robustness [6, 7, 8]. These approaches are orthogonal to our goal: they refine predictive distributions but do not specify the minimal prompt length for stability [2].
Mechanistic and empirical perspectives.
Mechanistic studies connect internal representations to emergent ICL behavior [9, 10]. Empirical benchmarks and toolkits provide standardized pipelines and evaluation practices [14, 16, 11]. Independent work observes that empirical knee-points often appear at substantially fewer examples than naive rules suggest [1], a phenomenon our theory explains through spectral coverage [17, 11].
Overall, our analysis complements these lines by supplying a lower-bound perspective that is both non-asymptotic and computable [18, 12].
3 Setup and Notation (re-stated for convenience)
Assumptions
  • (A1) Nominal assumption (strong). A deterministic feature map $\phi(x)$ is fixed and each $\phi(x)$ is sub-Gaussian with zero mean and covariance $\Sigma$. This assumption is adopted for the main theorem, under which all proofs are rigorous.
  • (A2) Approximate stationarity. In practice, the representation may depend on context index $t$. We model $\phi_{t}(x)=\phi(x)+\Delta_{t}(x)$ with $\|\Delta_{t}(x)\|_{2}\leq\varepsilon$ (uniformly in $x,t$ or in expectation). When $\varepsilon$ is small, our lower bound degrades smoothly by an additive correction term $g(\varepsilon)$ (see Appendix B).
  • (A4) Weak dependence. When the context induces dependence among features $\{\phi_{t}(x)\}$, we assume a mixing or martingale-difference condition. In this case, matrix concentration inequalities for dependent sequences apply, yielding bounds of the same form up to a dependence factor (Appendix C).
All formal proofs in the main text are carried out under the nominal assumption (A1). The relaxed settings (A2)–(A4) are analyzed in Appendix D, where we show that the main results remain valid up to controlled perturbation terms.
  • A frozen feature map $\phi:\mathcal{X}\to\mathbb{R}^{d}$ is assumed, satisfying $\mathbb{E}_{x\sim\mathcal{D}}\phi(x)=0$ and $\|\phi(x)\|_{\psi_{2}}\leq\sigma$ (the vector $\psi_{2}$-Gaussian norm).
  • The covariance is $\Sigma=\mathbb{E}[\phi\phi^{\top}]\in\mathbb{R}^{d\times d}$, with rank $r=\operatorname{rank}(\Sigma)$ (if $r<d$ we discuss on the support subspace). Denote the smallest nonzero eigenvalue by $\lambda_{r}(\Sigma)>0$ and for brevity write $\lambda_{\min}(\Sigma)=\lambda_{r}(\Sigma)$.
    $H_{K}=\begin{bmatrix}\phi(x_{1})\\ \vdots\\ \phi(x_{K})\end{bmatrix}\in\mathbb{R}^{K\times d},\qquad\widehat{\Sigma}_{K}=\frac{1}{K}H_{K}H_{K}^{\top}=\frac{1}{K}\sum_{k=1}^{K}X_{k},$
    where $X_{k}=\phi(x_{k})\phi(x_{k})^{\top}$ is a PSD matrix.
  • Goal: give the minimal $K$ (parameterized by $\Sigma,\sigma,r$, or provide an observable two-stage estimator) such that
    $\Pr\big(\lambda_{\min}(\widehat{\Sigma}_{K})\geq\delta\big)\geq 1-\xi,$
    where $0<\delta<\lambda_{\min}(\Sigma)$ and $\xi\in(0,1)$.
3.1 Basic matrix tail bound (a one-sided matrix Bernstein for the minimum eigenvalue)
Define centered matrices $Y_{k}:=X_{k}-\Sigma$, so $\mathbb{E}Y_{k}=0$ and
$\widehat{\Sigma}_{K}-\Sigma=\frac{1}{K}\sum_{k=1}^{K}Y_{k}.$
We need to control $\lambda_{\min}(\widehat{\Sigma}_{K})=\lambda_{\min}\!\Big(\Sigma+\frac{1}{K}\sum_{k=1}^{K}Y_{k}\Big)$. By a matrix Bernstein inequality (or a lower-tail form of Tropp-style matrix concentration), there exist constants $c,C>0$ such that for any $t\geq 0$,
where
Set $t=K\Delta$ where $\Delta:=\lambda_{\min}(\Sigma)-\delta>0$. Substituting into (1) and adjusting scaling yields for any $K$,
$\Pr\big(\lambda_{\min}(\widehat{\Sigma}_{K})\leq\delta\big)\leq r_{\mathrm{eff}}\exp\!\Big(-\frac{K\Delta^{2}}{2\widetilde{v}+\tfrac{2}{3}R\Delta}\Big),$ (2)
3.2 Controlling parameters under the sub-Gaussian assumption (truncation and fourth-moment bounds)
3.2.1 Fourth-moment bound
$\mathbb{E}\|\phi\|^{4}\leq C_{1}\sigma^{4}.$ (3)
Hence
Using the operator inequality $(X-\Sigma)^{2}\preceq 2X^{2}+2\Sigma^{2}$ we get
$\widetilde{v}=\big\|\mathbb{E}[(X-\Sigma)^{2}]\big\|\leq 2\mathbb{E}\|X\|^{2}+2\|\Sigma\|^{2}\leq 2C_{1}\sigma^{4}+2\|\Sigma\|^{2}.$ (5)
3.2.2 Truncation to control R
Typically $\|X_{k}\|=\|\phi(x_{k})\|^{2}$ is not almost surely bounded, but by the sub-Gaussian tail (there exists $c_{1}>0$)
$\Pr(\|\phi\|\geq t)\leq 2\exp\!\Big(-\frac{c_{1}t^{2}}{\sigma^{2}}\Big).$
Choose threshold $t_{0}=C_{2}\sigma\sqrt{\log(8K/\xi)}$ (constant $C_{2}$ depends on $C_{1}$, with $C_{1}C_{2}^{2}=1$). Then for a single sample
$\Pr(\|\phi\|>t_{0})\leq\frac{\xi}{4K}.$
By a union bound over the $K$ samples, with probability at least $1-\xi/4$,
On this high-probability event we may take
3.3 Substitute into (2) to obtain a sample-size lower bound (one main result)
Let $\Delta=\lambda_{\min}(\Sigma)-\delta>0$. If
$\Pr(\lambda_{\min}(\widehat{\Sigma}_{K})\geq\delta)\geq 1-\xi.$
3.4 A more concise, commonly used operator-norm concentration (avoid truncation and the self-consistency)
Using (9) to ensure $\lambda_{\min}(\widehat{\Sigma}_{K})\geq\delta$ (appendix explains why this suffices) we require
$C\|\Sigma\|\Big(\sqrt{\frac{r_{\mathrm{eff}}\log(2/\xi)}{K}}+\frac{r_{\mathrm{eff}}\log(2/\xi)}{K}\Big)\leq\Delta.$
$K\geq C^{\prime}\frac{\|\Sigma\|^{2}r_{\mathrm{eff}}\log(2/\xi)}{\Delta^{2}},$ (10)
3.5 A two-stage observable (plug-in) scheme and finite-sample guarantee
Two-stage algorithm (high level)
  1. 1.
    $\widehat{\lambda}_{0}:=\lambda_{\min}(\widehat{\Sigma}_{K_{0}}),\qquad\|\Sigma\|_{\wedge}:=\|\widehat{\Sigma}_{K_{0}}\|,\qquad\widehat{r}_{\mathrm{eff}}:=\frac{\operatorname{tr}(\widehat{\Sigma}_{K_{0}})}{\|\widehat{\Sigma}_{K_{0}}\|}.$
  2. 2.
    $\lambda:=\widehat{\lambda}_{0}-C\|\Sigma\|_{\wedge}\Big(\sqrt{\frac{\widehat{r}_{\mathrm{eff}}\log(4/\xi)}{K_{0}}}+\frac{\widehat{r}_{\mathrm{eff}}\log(4/\xi)}{K_{0}}\Big).$ (11)
    Here the constant $C$ is the same as in (9), and the failure probability is allocated to $\xi/2$ (details below).
  3. 3.
    $K\geq C^{\prime}\frac{\|\Sigma\|_{\wedge}^{2}\widehat{r}_{\mathrm{eff}}\log(2/\xi)}{\widehat{\Delta}^{2}}.$
  4. 4.
    The final total sample size is $K_{0}+K$, and with total failure probability $\leq\xi$ one guarantees $\lambda_{\min}(\widehat{\Sigma}_{K})\geq\delta$.
Finite-sample correctness (theoretical guarantee)
Using (9) for the first stage: if $K_{0}$ satisfies
then with probability at least $1-\xi/2$,
$|\widehat{\lambda}_{0}-\lambda_{\min}(\Sigma)|\leq\frac{\lambda_{\min}(\Sigma)}{4},\qquad\big|\|\widehat{\Sigma}_{K_{0}}\|-\|\Sigma\|\big|\leq\frac{\|\Sigma\|}{4}.$
Hence (using simple algebra aa/4=3a/4a-a/4=3a/4, etc.)
$\lambda\geq\lambda_{\min}(\Sigma)-\frac{\lambda_{\min}(\Sigma)}{4}-C\cdot\frac{5\|\Sigma\|}{4}\Big(\sqrt{\frac{r_{\mathrm{eff}}\log(4/\xi)}{K_{0}}}+\frac{r_{\mathrm{eff}}\log(4/\xi)}{K_{0}}\Big)\geq\frac{7\lambda_{\min}(\Sigma)}{16},$

### v2

@@ Abstract
[…]
Machine Learning, ICML
Figure 1: Schematic comparison of ICL prompt-length selection. (a) Conventional trial-and-error or heuristic approaches leave uncertainty about the minimal effective number of demonstrations (K). (b) Our proposed spectral-coverage method provides a theoretically grounded, observable proxy to determine a stable minimum prompt length (K*), ensuring reliable in-context learning.
@@ 1 Introduction
In-context learning (ICL) has become a key paradigm for large language models to adapt to downstream tasks, as it enables task inference without parameter updates, solely through prompting(Wei et al., 2022; Tanwar et al., 2023; Dong et al., 2024). However, the output of ICL is sensitive to the selection and order of specific examples in the prompt(Xie et al., 2022; Zhao et al., 2021; Chen et al., 2023; Li and Qiu, 2023; Fang et al., 2025), making its stability difficult to guarantee. An unresolved core question is: how many examples are needed to ensure stable ICL predictions? The lack of a computable answer to this question leads to reliance on heuristics or excessively long prompts in practice, sacrificing efficiency and failing to provide reliability guarantees.
This paper studies the stability of ICL and the minimal necessary prompt length. We formally define ICL stability in a probabilistic sense: for the same query, two independently sampled sets of examples should induce similar model output distributions. Directly evaluating this stability requires extensive repeated queries to the model, which is costly. To address this, we introduce an observable proxy objective: spectral coverage of example representations, specifically manifested as the minimum eigenvalue of the regularized empirical second-moment (uncentered covariance) matrix. Our goal is to estimate, from data, the minimum number of examples $K^{*}$ required to meet this proxy objective under given stability tolerances $(\tau,\xi)$, and to provide non-asymptotic theoretical guarantees.
In black-box API settings, the generator typically does not expose hidden states. We therefore instantiate the representation map $\phi$ with a frozen external encoder $e(\cdot)$ and compute all spectral quantities in this encoder space. Our finite-sample guarantees are for this observable proxy (Remark 3.2); empirically, we assess proxy–generator alignment via direct resampling-based stability estimates (Appendix L) and a validation-only calibration step (Section 3.5).
Previous work has analyzed ICL from various perspectives: theoretical studies have explored learnability under mixture models(Chan et al., 2024), sample complexity of Transformers(Bai et al., 2023), and asymptotic learning curves(Jeon et al., 2024); calibration methods have improved confidence in few-shot predictions(Abbas et al., 2024; Cho et al., 2025; Tan et al., 2025); mechanistic research has linked internal representations to ICL behavior(Dai et al., 2023; Hendel et al., 2023). However, none of these works provides a non-asymptotic, computable lower bound for estimating the minimum number of demonstrations required for stable ICL in practical tasks. Existing rules of thumb are often overly conservative and cannot establish verifiable consistency with the actual behavior of models (e.g., empirical knee-points).
To this end, we propose a framework based on spectral coverage theory to estimate the minimal prompt length required for stable ICL. Our central insight is that the spectral lower bound of the empirical second-moment matrix of example representations provides an effective and computable proxy for ICL stability. Specifically, under a linearized ridge-regression view of in-context learning, sufficient spectral coverage controls the conditioning of the induced inverse geometry, thereby limiting the sensitivity of predictions to perturbations in the demonstration set. Building on this insight, we design an observable two-stage algorithm: the first stage uses a small initial sample to estimate spectral statistics and construct a confidence lower bound, while the second stage computes the additional sample size needed to achieve a target level of spectral coverage. Crucially, this procedure preserves non-asymptotic theoretical guarantees without requiring prior knowledge of the population second-moment matrix.
We conduct extensive experiments on multiple tasks covering text classification, multi-step reasoning, and retrieval-augmented generation, involving several commercial API models. Our theory yields an explicit sufficient sample-size requirement $K\geq K_{\mathrm{th}}$ for the spectral-coverage proxy event (and hence reduced sensitivity to demonstration perturbations in the stylized view). Empirically, the resulting estimates $K_{\mathrm{th}}$ also form a conservative upper envelope for observed accuracy knee-points $K_{\mathrm{knee}}$ under fixed decoding protocols. The results show that raw error ratio $\approx 1.3-2.5\times$, defined as the ratio between the predicted prompt length and the empirical knee-point). Through a lightweight calibration step using only a validation set (learning a global scaling factor and using quantile eigenvalues), the gap can be significantly tightened (calibrated error ratio $\approx 1.03-1.20\times$). We discuss robustness extensions of the assumptions in the appendix for representation drift, heavy-tailed features, and weak dependence.
The contributions of this paper are as follows:
Theory (sufficient condition & threshold interpretation): We derive a non-asymptotic sufficient condition for spectral coverage under sub-Gaussian representations, expressed as an explicit lower bound on the required prompt length $K$. Interpreting this condition yields a conservative upper bound on the (unknown) stability threshold, and we further analyze robustness under representation drift, heavy tails, and weak dependence. Observable estimator (no population priors): We design a fully observable two-stage procedure that estimates the required spectral statistics from a small pilot prompt and returns a concrete prompt length with a prescribed failure probability.
Empirical alignment (knee, stability diagnostics, calibration): Across datasets, task families, and commercial API models, we verify that the predicted thresholds upper-bound empirical knee-points, and that a lightweight validation-only calibration (global scale and quantile level) substantially tightens the gap while preserving conservative ordering.
@@ 2 Related Works
Our work on the minimal prompt length for stable ICL sits at the intersection of three research strands: theoretical analysis of ICL, empirical studies of its reliability, and mechanistic explanations of its emergence. We briefly review these directions and highlight the gap we fill: the lack of a theoretically grounded, practically observable, non-asymptotic, and computable lower bound for ICL stability.
Theoretical foundations of ICL. A growing body of work formalizes ICL within established learning frameworks. PAC-style analyses establish learnability conditions under mixture models, while other studies treat Transformers as algorithms with task-dependent sample complexity(Frei and Vardi, 2024; Magen and Vardi, 2025; Li et al., 2025). Asymptotic analyses, particularly in linearized settings, describe phase transitions in learning curves, and task-specific results for regression and structured prediction further refine these analyses(Chang et al., 2025; Takanami et al., 2025). While these provide key asymptotic insights, they often lack non-asymptotic, quantitative guarantees that prescribe a concrete number of demonstrations $K$ for a given task and model—a critical need for practical prompt design. Our work fills this gap by deriving a non-asymptotic lower bound on $K$ based on spectral coverage of demonstration representations, offering a certificate of stability beyond mere eventual convergence.
Calibration and reliability in few-shot learning. Another line of research focuses on calibration methods to improve trustworthiness in ICL predictions. These techniques typically adjust output distributions to enhance predictive confidence and robustness(Gundem et al., 2025; Shen et al., 2025). However, they are orthogonal to our core goal: calibration methods generally assume a fixed prompt and aim to improve the interpretability of its outputs; our goal is to determine the minimal prompt length needed for the output distribution itself to remain stable under demonstration resampling. Thus, our spectral coverage proxy addresses a prerequisite question not tackled by calibration.
Mechanistic and empirical perspectives. Understanding how ICL emerges internally has been explored through mechanistic studies linking Transformer internal representations to ICL behavior(Li and Qiu, 2023; Fang et al., 2025). Empirically, comprehensive benchmarks and toolkits have standardized evaluation pipelines(Chen et al., 2025; Rieff et al., 2025). Notably, independent empirical work observes that performance knee-points often require far fewer demonstrations than heuristic rules suggest(Liu et al., 2022; Hendel et al., 2023). Our theory provides a direct mechanistic and quantitative explanation for this phenomenon: knee-points coincide with the moment when the empirical second-moment (uncentered covariance) matrix of demonstrations achieves sufficient spectral coverage, a condition we formalize and make computable. While these empirical and mechanistic studies document phenomena and mechanisms, our analysis provides a predictive why and when from a statistical learning perspective.
Our positioning. In summary, prior theoretical work offers asymptotic characterizations rather than practical lower bounds; calibration research optimizes outputs for fixed prompts rather than determining prompt size; empirical/mechanistic studies describe phenomena that we can now quantitatively explain. Our contribution complements these lines by introducing a spectral coverage proxy for ICL stability, deriving its non-asymptotic lower bound, and providing an efficient two-stage algorithm to estimate it from data. This approach offers a missing normative, sample-complexity-oriented perspective, turning qualitative observations about knee-points into a quantitatively verifiable and actionable stability condition.
3 Theory
3.1 Problem setup and proxy objective
A $K$-shot prompt is formed from $K$ demonstrations $S_{K}:=\{(x_{i},y_{i})\}_{i=1}^{K}$ and a query input $x_{q}$ via a fixed template $\mathrm{Prompt}(\cdot)$. Given $\mathrm{Prompt}(S_{K},x_{q})$, a (black-box) language model induces an output distribution $p_{\theta}(\cdot\mid\mathrm{Prompt}(S_{K},x_{q}))$ (e.g., over next tokens, or over labels under a verbalizer). The randomness we study comes from resampling (and, when relevant, re-ordering) the demonstrations in the prompt.
Definition 3.1 (ICL stability at $(\tau,\xi)$).
Fix a query input $x_{q}$ and a prompting template. Let $S_{K}$ and $S^{\prime}_{K}$ be two independent sets of $K$ demonstrations drawn i.i.d. from the same task distribution. Define $P_{K}:=\mathrm{Prompt}(S_{K},x_{q})$ and $P^{\prime}_{K}:=\mathrm{Prompt}(S^{\prime}_{K},x_{q})$. ICL is stable at length $K$ if
$\Pr\!\Big(d\big(p_{\theta}(\cdot\mid P_{K}),\;p_{\theta}(\cdot\mid P^{\prime}_{K})\big)\leq\tau\Big)\;\geq\;1-\xi.$
where $d(\cdot,\cdot)$ is a divergence between distributions (e.g., total variation, KL, or Jensen–Shannon).
Intuition and example.
Definition A.1 formalizes stability in terms of the agreement between output distributions induced by two independently sampled prompts of the same length. Concretely, consider a classification task where each prompt consists of $K$ labeled demonstrations. Two prompts $P_{K}$ and $P^{\prime}_{K}$ are formed by independently sampling and ordering demonstrations from the same task pool. If the model is unstable at length $K$, small changes in the selected demonstrations may lead to noticeably different output distributions—for example, flipping the predicted label or substantially redistributing probability mass among competing classes. In contrast, stability implies that, despite variation in the demonstrations, the model produces highly similar predictive distributions, reflecting insensitivity to the specific realization of the prompt. This notion captures robustness to prompt composition rather than average accuracy, and is task-agnostic.
Remark 3.2 (Scope of the proxy and instantiation of $\phi$).
Definition A.1 is stated for a general LLM output distribution. Our theory does not model the full generator; instead, it analyzes a computable spectral-coverage proxy under a frozen representation map $\phi$ and standard concentration assumptions. In black-box API settings, the generator typically does not expose hidden states, so in practice $\phi$ can be instantiated using a fixed external encoder. In our experiments, $\phi(x)$ is instantiated as a fixed sentence-level embedding $e(x)$ of the input text, and we form $\widehat{\Sigma}_{K}:=\frac{1}{K}\sum_{i=1}^{K}e(x_{i})e(x_{i})^{\top}$ from the selected demonstrations, denoting the empirical second-moment (uncentered covariance) matrix of the representation vectors induced by the $K$ demonstrations. (see Appendix C for encoder and preprocessing details). All finite-sample guarantees in this section concern the resulting proxy quantities computed in the $\phi$-space.
Directly verifying Definition A.1 is expensive because it requires repeated model queries under many resampled prompts. We therefore study an observable sufficient condition based on the spectral coverage of demonstration representations. Let $\phi:\mathcal{X}\to\mathbb{R}^{d}$ be a fixed feature map and define the (uncentered) population second-moment matrix
$\Sigma:=\mathbb{E}\!\left[\phi(x)\phi(x)^{\top}\right]\succeq 0,$
which does not require $\mathbb{E}[\phi(x)]=0$. Given i.i.d. demonstrations $x_{1},\ldots,x_{K}$, define
$H_{K}$ $:=\begin{bmatrix}\phi(x_{1})^{\top}\\ \vdots\\ \phi(x_{K})^{\top}\end{bmatrix}\in\mathbb{R}^{K\times d},$ (1)
$\widehat{\Sigma}_{K}$ $:=\frac{1}{K}H_{K}^{\top}H_{K}=\frac{1}{K}\sum_{k=1}^{K}X_{k},\;\;\;X_{k}:=\phi(x_{k})\phi(x_{k})^{\top}.$
For numerical robustness we use ridge regularization: for a ridge level $\rho\geq 0$, let
$\widehat{\Sigma}_{K,\rho}:=\widehat{\Sigma}_{K}+\rho I_{d}.$
Our primary proxy objective is to find the minimal $K^{\star}$ such that
$\Pr\!\big(\lambda_{\min}(\widehat{\Sigma}_{K,\rho})\geq\delta\big)\;\geq\;1-\xi,$ (2)
for a target floor $\delta>\rho$ and failure tolerance $\xi\in(0,1)$. This proxy is equivalent to requiring $\lambda_{\min}(\widehat{\Sigma}_{K})\geq\delta-\rho$ but is more stable numerically and aligns with ridge/regularized linearized views of ICL.
To reduce sensitivity to the extreme tail of the spectrum, we will also consider a quantile surrogate. Let $\lambda^{(1)}(A)\leq\cdots\leq\lambda^{(d)}(A)$ denote the eigenvalues of a PSD matrix $A$ in nondecreasing order, and define
$\lambda_{q}(A):=\lambda^{(\lceil qd\rceil)}(A),\qquad q\in(0,1).$
In practice we often use $\lambda_{q}(\widehat{\Sigma}_{K,\rho})$ (e.g., $q\in[0.05,0.2]$) to obtain less conservative yet still monotone proxies. Throughout the main analysis we adopt the following standard nominal condition on $\phi(x)$.
(A1)Sub-Gaussian features.The feature vector $\phi(x)\in\mathbb{R}^{d}$ is sub-Gaussian (e.g., $\|\phi(x)\|_{\psi_{2}}\leq\sigma$) and $\Sigma=\mathbb{E}[\phi(x)\phi(x)^{\top}]$ exists.
This assumption should be viewed as a working condition enabling clean non-asymptotic analysis; in the appendix, we discuss extensions to heavy-tailed features, representation drift, weak dependence, and empirically evaluate robustness beyond this idealized setting.
We discuss robustness extensions under representation drift, heavier tails, and weak dependence in the appendix (moved from the original main-text derivations).
Notation. Let $r=\mathrm{rank}(\Sigma)$ and denote the smallest nonzero eigenvalue by $\lambda_{r}(\Sigma)>0$; for brevity we write $\lambda_{\min}(\Sigma)=\lambda_{r}(\Sigma)$. We use $\|\cdot\|$ for the operator norm and $\mathrm{tr}(\cdot)$ for trace. “Effective rank” will be measured by
$r_{\mathrm{eff}}:=\frac{\mathrm{tr}(\Sigma)}{\|\Sigma\|}\leq r,$
which naturally appears in matrix concentration bounds.
3.2 Sample complexity bound
We study the proxy objective in (2) and derive a non-asymptotic sample complexity guaranteeing a spectral “floor” for the ridge-regularized empirical second moment. The full derivation (matrix Bernstein lower tail, truncation, and variance proxy control) is moved to the appendix; here we state only the final, implementation-facing bound.
Define the spectral margin
$\Delta_{\rho}\;:=\;\lambda_{\min}(\Sigma)+\rho-\delta.$ (3)
The regime of interest is $\Delta_{\rho}>0$, i.e., the target floor $\delta$ is below the population ridge floor $\lambda_{\min}(\Sigma)+\rho$.
Proposition 3.3 (Spectral coverage from $K$ demonstrations).
Assume (A1) (sub-Gaussian features) and let $\xi\in(0,1)$. There exist universal constants $c,C>0$ such that if $\Delta_{\rho}>0$ and
$K\;\geq\;C\cdot\frac{\|\Sigma\|^{2}}{\Delta_{\rho}^{2}}\,r_{\mathrm{eff}}\,\log\!\Big(\frac{c\,r_{\mathrm{eff}}}{\xi}\Big),$ (4)
then the proxy objective holds:
$\Pr\!\big(\lambda_{\min}(\widehat{\Sigma}_{K,\rho})\geq\delta\big)\;\geq\;1-\xi.$ (5)
The leading dependence in (4) is the familiar
$K\;\asymp\;\frac{\text{(scale)}^{2}\times\text{(effective dimension)}\times\log(1/\xi)}{\text{(spectral margin)}^{2}},$
where the scale is $\|\Sigma\|$, the effective dimension is $r_{\mathrm{eff}}=\mathrm{tr}(\Sigma)/\|\Sigma\|$, and the margin is $\Delta_{\rho}$. Intuitively, larger feature variance (larger $\|\Sigma\|$) and higher intrinsic dimensionality (larger $r_{\mathrm{eff}}$) require more demonstrations to ensure that the empirical covariance is well-conditioned along all directions, while a larger margin $\Delta_{\rho}$ makes the guarantee easier to satisfy.
The appendix also provides an explicit non-asymptotic sufficient condition with additional higher-order and logarithmic terms (arising from controlling per-sample operator norms and a self-consistency step). We omit that longer expression in the main text to keep the theory readable; see the appendix for the complete chain and constants.
Although Proposition 3.3 is stated in terms of population quantities, it is designed to be used via plug-in estimates of $\|\Sigma\|$, $r_{\mathrm{eff}}$, and $\lambda_{\min}(\Sigma)$. Section 3.3 introduces an observable two-stage procedure that estimates these terms from data and returns a concrete prompt length $K$ with the desired failure probability.
3.3 An observable two-stage estimator
Proposition 3.3 suggests choosing $K$ using population quantities $\lambda_{\min}(\Sigma)$, $\|\Sigma\|$, and $r_{\mathrm{eff}}$. In practice these are unknown, and we seek an observable procedure that (i) estimates the needed scale and intrinsic dimension from data, and (ii) returns a concrete prompt length with a prescribed failure probability.
We use a small pilot prompt to obtain a coarse but reliable estimate of the spectrum, and then plug these estimates into the readable bound (4). To guard against optimism from finite-sample noise, we include a safety margin on the spectral gap (3) and allocate the total failure probability across the two stages.
Given $K_{0}$ demonstrations, form $\widehat{\Sigma}_{K_{0}}$ and $\widehat{\Sigma}_{K_{0},\rho}=\widehat{\Sigma}_{K_{0}}+\rho I$ as in Section 3.1, and compute the observable quantities
$\widehat{\lambda}_{0}:=\lambda_{\min}(\widehat{\Sigma}_{K_{0}}),\widehat{s}_{0}:=\|\widehat{\Sigma}_{K_{0}}\|,\widehat{r}_{\mathrm{eff},0}:=\frac{\mathrm{tr}(\widehat{\Sigma}_{K_{0}})}{\|\widehat{\Sigma}_{K_{0}}\|}.$ (6)
We then define a conservative estimated margin
$\widehat{\Delta}_{\rho}\;:=\;\big(\widehat{\lambda}_{0}-\eta\big)+\rho-\delta,$ (7)
where $\eta\geq 0$ is a user-chosen slack (or a bound-driven correction; see appendix for one principled choice). If $\widehat{\Delta}_{\rho}\leq 0$, we increase $K_{0}$ and repeat.
Once $\widehat{\Delta}_{\rho}>0$, we output a plug-in length
$\widehat{K}\;:=\;\left\lceil C\cdot\frac{\widehat{s}_{0}^{\,2}}{\widehat{\Delta}_{\rho}^{\,2}}\,\widehat{r}_{\mathrm{eff},0}\,\log\!\Big(\frac{c\,\widehat{r}_{\mathrm{eff},0}}{\xi_{2}}\Big)\right\rceil,$ (8)
where $\xi_{2}$ is the failure budget assigned to Stage II and $c,C$ are the same universal constants as in Proposition 3.3. In experiments we optionally apply a single global calibration factor to reduce conservatism (Section on calibration).